solve each proportion
2/5 = x/40
15/3 = x-3/5
x+1/3 = 7/2
3x-5/4 = -5/7
x-2/4 = x+4/2
x/1.12 = 1/5

Answer by checkley79(3341)

2/5 = x/40 CROSS MULTIPLY.
5X=2*40
5X=80
X=80/5
X=16 ANS.

15/3 = (x-3)/5
3(X-3)=5*15
3X-9=75
3X=75+9
3X=84
X=84/3
X=28 ANS.
##################################################
I'LL LEAVE THE REST FOR YOU TO PRACTICE WITH.
x+1/3 = 7/2
3x-5/4 = -5/7